UVernmen Alternatively, by how much would the government have to change ta the fall in planned investment t have to change its purchases to re ly to full employment? pose that the governments budget is initially in balance, with ment spending equal to taxes colle government ers could do to not want to violate the balanced-budget law? d Sup cted. A bal nced-budget law for from running a deficit. Is there anything that fiscal pol restore full employment in this economy, assuming 7. An economy is described by the following equations: C 40 + 0.8(Y-T) IP 70 G 120 NX 10 T 150 Y 580
media%2F5ff%2F5ffcc0e3-36da-449a-afd2-6c
Answer #1

a) PAE = Y

C + I + G + NX = Y

40 + 0.8(Y-150) + 70 + 120 + 10 = Y

120 + 0.8Y = Y

This is the required equation.

b) Table is given below. The equilibrium level of output is 600.

Y (Output) PAE
0 120
100 200
200 280
300 360
400 440
500 520
600 600
700 680

c) Full employment output is 580 and current equilibrium is 600. Hence inflationary gap is 20. Government spending should be reduced by Gap/spending multiplier or taxes should be increased by Gap/tax multiplier

Spending multiplier is 1/1-mpc = 5 and tax multiplier = -mpc/1-mpc = -4

Hence Government spending should be reduced by Gap/spending multiplier = 20/5 = $4 or taxes should be increased by Gap/tax multiplier = 20/4 = $5.

Actual expenditure Expenditure Planned PE1 MPCYAT expenditure 0.8 x 5 PE2 4Y-20 45 y, =580 y, -600 Income, output, Y 2

d) Full employment output is 630 and current equilibrium is 600. Hence recessionary gap is 30. Government spending should be increasd by Gap/spending multiplier or taxes should be decreased by Gap/tax multiplier

Hence Government spending should be increased by Gap/spending multiplier = 30/5 = $6 or taxes should be decreased by Gap/tax multiplier = 30/4 = $7.5

Expenditure Actual expenditure Planned ΔG_S6 expenditure 45 Y600 Y -630 Income, output, Y 2